OpenStudy (anonymous):

Evaluate\[\lim_{x \rightarrow \infty} \text{Arcsin}(\frac{x+2}{x+1})\]

hartnn (hartnn):

arcsin(1)

OpenStudy (anonymous):

pi/2?

hartnn (hartnn):

or we can't take lim inside??

OpenStudy (anonymous):

we can

OpenStudy (anonymous):

so what is it?

hartnn (hartnn):

pi/2+2npi

OpenStudy (anonymous):

@mukushla...i got pi/2. i took tke limit inside; there i divided each element by the variable with the highest degree which is obviously x. i had limit as x app infinity (1+2/y)/(1+1/y) where y is a very large number. 2/y and 1/y then approach 0 rendering everything in the bracket all equals 1. i then took the arcsine of 1

OpenStudy (anonymous):

u guys still missing something

hartnn (hartnn):

x cannot take the value of infinity,arcsin is bounded.

OpenStudy (anonymous):

so limit does not exist

hartnn (hartnn):

u sure we can take lim inside?

OpenStudy (anonymous):

yes like this\[\lim_{x \rightarrow \infty} \text{Arcsin}(\frac{x}{x+1})=\text{Arcsin} (1)=\frac{\pi}{2}\]

OpenStudy (anonymous):

hey man u lost me...what did u do to whats inside the bracket?